# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 3.3 Physics, chemistry, and mechanics helping life science applications
The RD systems have also been applied to high energy physics (HEP) and statistical physics. RD models proposed by Balitsky and by Jalilian-Marian, Iancu, McLerran, Leonidov, Kovner, and Weigert (JIMLKW) have been used to understand the high energy hard scattering near or at the unitarity limit balitsky1996 ; fukushima2006 . Ikeda, Nagasawa, and Watanabe have introduced the branching Brownian motion in statistical physics ikeda1965 ; ikeda1968 , which can be associated with the classical Fisher-KPP equation, as shown by McKean in mckean1975 . Nonlocal interactions can occur as a result of long-range forces such as gravitational or electromagnetic interactions and can have a considerable impact on the behaviour and dynamics of multi-body systems. These interactions are critical in astrophysics for analyzing the dynamics of galaxies because gravitational forces between stars and other celestial entities reach over large distances binney1987 . In galaxies, nonlocal interactions between stars and dark matter particles regulate galactic systems’ overall structure, rotation, and stability. Furthermore, the authors of spergel2000 examined the self-interacting dark matter in cosmology, emphasizing the importance of nonlocal interactions in the dynamics of multi-body systems on cosmological scales.
Nonlocal variational formulations have been used to derive and study nonlocal PDEs, which appear in diverse physical and biological systems. For example, the nonlocal aggregation equation (42) has been adapted to describe biological organisms bernoff2016 ; bertozzi2007 ; bertozzi2012 , and it minimizes energy
@@
and it can be used to analyze the existence and stability of equilibrium points bernoff2011 ; carrillo2014 ; simione2015 . Despite its popularity, it is not biological for a minimizer of mass $M$; doubling the mass yields a double-density minimizer. On the other hand, the dimensionless aggregation model (43) does form well-spaced groups. This model minimizes the energy topaz2006 ; bedrossian2011 ; burger2014 :
@@
These nonlocal variational problems are also used in image denoising and restoration tasks. The nonlocal operators capture the similarities between different image patches and help preserve important structures while reducing noise gilboa2009 . Furthermore, these problems are employed to study material’s deformation and fracture behaviour with long-range interactions and discontinuities, such as brittle materials and nanomaterials.
Nonlocal models also describe Ohmic heat production in different industrial processes, e.g., food sterilization, thermistor devices kavallaris2018 , etc. Using Ohmic heating, the nonlocal model of food sterilization in the one-dimensional spatial domain $(0,1)$ is given as kavallaris2002 :
@@
with the initial condition $u(x,0)=u_{0}(x),0<x<1$ and the boundary condition $u(0,t)=u_{x}(1,t)=0,t>0$. Here, the function $f>0$ is the dimensionless electrical reactivity, and $\lambda>0$ is the square of the applied potential difference.
@@
Nonlocal models are used in many domains, including mechanics, materials science, and continuum physics. Nowadays, nanotechnology is one of the main technological interests in the research community, and nonlocal continuum mechanics could play a potential role in analyzing their applications peddieson2003 . Piezoelectric materials are used to make ultrasound scanners, and these sensors use the converse piezoelectric effect to send out sound vibrations reflected by muscles, organs, etc. manjon2018 . These ultrasound devices also use the piezoelectric effect to detect the reflected sound. Then, they are displayed as images so that doctors can see those highly detailed 3D images for treatments. Other types of applications of these materials in the medical field are the making of sensors for electronic stethoscopes, hearing aids, flow meters to monitor vascular health and both external and internal sensors for health status monitoring. These materials are also used in the automobile industry, e.g., knock sensors, tire pressure monitors, airbag and seatbelt sensors, gyroscopes, accelerometers, and engine fluid detection varghese2015 . Piezoelectric materials and devices are increasingly important in the digital and physical world.
@@
Quantum dots are examples of a low-dimensional nanostructure, and researchers have been focused on these due to their several nanoelectronic applications zhang2005 ; zhang2005a . In the context of nano-inclusions and quantum dots, the size dependency of mechanical strain at the nanoscale has been brought to light in recent publications. Piezoelectricity is the growing importance of nonlocal models in the context of smart materials and structure yang2006 . Gradient theories are closer to microscopic theories like lattice dynamics than classical continuum theories. They are still applicable when the characteristic length of a problem is so small that classical continuum theories begin to fail yang2006 . The application of nonlocal continuum theory to nanotechnology was initially addressed by Peddieson et al. peddieson2003 . This model has been extended to the buckling and vibration analysis of multiwall carbon nanotubes lu2006 .
@@
Peridynamics is a nonlocal theory and is one of the key methods for modelling biological material fracture madenci2014 . The main advantage of this nonlocal model is that it deals with continuous and discrete media lejeune2021 . It has been applied to find (i) the relationship between fracture and healing in cortical bone deng2009 ; (ii) fracture in anisotropic cortical bone ghajari2014 ; (iii) fracture in a complex wood microstructure madenci2014 ; (iv) rupture in a micron-scale biological membrane taylor2016 ; (v) fracture in a porous material chen2019 . Furthermore, peridynamics is used to model tissue growth and shrinkage. The distribution of division angle orientations influences tissue scale growth, and researchers have been trying to predict it in an individual cell gillies2011 . Authors in lejeune2021 proposed a mechanics-based model with the help of the peridynamics framework. In addition, the authors have shown cell death differences on cellular and larger length scales. In particular, cell death may create gaps between cells in some cases, leading to tissue shrinkage in other cases.
@@
The nonlocal effects in heat-mass transfer arise in a wide variety of modern practical applications in materials sciences, e.g., metallic nanowires sellitto2012 , high-power laser melting sobolev1994 , colloidal solidification sobolev2012 , frontal polymerization sobolev1998 , etc. The classical diffusion theory based on Fick’s law suggests the diffusion flux at a spatial point $x$, and a time moment $t$ depends on the concentration gradient at the same space-time point. These classical heat conduction laws apply at macro-temporal and spatial scales but not to thermal non-equilibrium situations, e.g., cryogenic engineering, laser-aided material processing, high-rate heat transfer in rarefied media, etc. Nowadays, researchers believe that heat-mass transport is an inherently nonlocal phenomenon. For instance, in one spatial dimension, a particle can come to a spatial point $x$ at a time moment $t+\tau$ from a spatial point $(x-l)$ or $(x+l)$ where it was at a time moment $t$. It implies a time lag $\tau$ between the diffusion flux and the particle concentration. This time lag $\tau$ is called relaxation time, particularly the mean collision time for gas particles. In the one-dimensional case, the net flux at a spatial point $x$ and time $t+\tau$ is given by
@@
After applying the zero-order approximation and Fick’s law, one can obtain the classical diffusion equation of the parabolic type. In addition, this approximation is valid for slow processes, i.e., $\tau\ll t_{0}$ and $l\ll h$, where $t_{0}$ and $h$ are the characteristic time and distances, respectively. For the other cases, e.g., $\tau\varpropto t_{0}$ and/or $l\varpropto h$, higher order approximation should be used as the nonlocal effects play a significant role. If the diffusion-like condition connects the small parameters:
@@
then, the first order approximation of (47) along with the mass conservation law, one can obtain sobolev1993 ; sobolev1997 ; sobolev2014 :
$\frac{\partial C}{\partial t}+\tau\frac{\partial^{2}C}{\partial t^{2}}=d\frac{\partial^{2}C}{\partial x^{2}}+d\frac{l^{2}}{3}\frac{\partial^{4}C}{\partial x^{4}}.$
This is a parabolic type nonlocal equation, and for the same limiting condition (48), a higher order approximation also gives the parabolic type equation. For the wavelike dynamics, the limiting relation between the small parameters $\tau$ and $l$ is
@@
$\frac{\partial C}{\partial t}+\tau\frac{\partial^{2}C}{\partial t^{2}}=d\frac{\partial^{2}C}{\partial x^{2}}.$
The limiting condition (49) also reduces to the hyperbolic type nonlocal equation for higher order approximation of (47).
@@ 4 Emerging Trends, Future Directions, and Outlook
The areas of mathematical modelling in biology are still evolving; various developing patterns in nonlocal mathematical models have been observed. As research progresses, these models are expected to play an increasingly significant role in addressing new challenges in the field of biology and life sciences. These trends reflect advances in computational approaches, enhanced experimental data availability, and a rising acknowledgement of the importance of nonlocal interactions in understanding complex biological phenomena. Some of the emerging trends in nonlocal mathematical models are the following.
@@
Multi-scale modelling: Biological systems often span multiple spatial and temporal scales, and nonlocal models provide a robust and powerful framework to bridge these scales and capture the intricate interactions between different components of the system, from molecular to cellular to tissue and organism levels, however, traditional modelling approaches often focus only on specific scales nieminen2002 ; shaat2020 .
@@
Human-environment systems, biodiversity, and ecological modelling: Human-environment systems represent the intricate relationship between human societies and their surrounding environments millennium2005 . In addition, biodiversity is the variety of life on Earth, encompassing different species, ecosystems, and genetic diversity. It is critical for maintaining ecological balance, delivering ecosystem services, and contributing to ecosystem resilience in the face of environmental change. Ecological modelling is developing and analyzing mathematical and computer models to replicate and comprehend the dynamics of ecosystems. Studying the spatio-temporal trends of human activities, such as deforestation rates, carbon emissions, and resource consumption, provides insights into the evolving nature of anthropogenic impacts on the environment. These models assist researchers in predicting the impact of human activities on biodiversity, studying ecosystem responses, and developing sustainable resource management techniques.
@@
Collective cell migration: Collective cell migration is a phenomenon where groups of cells move together to achieve specific functions, such as embryonic development, wound healing, or cancer invasion sherratt1990 ; maini2004 ; chen2020 . These nonlocal interactions can influence cell behaviour, signalling, and collective dynamics in complex biological systems. These interactions play a significant role in coordinating the movements of cells within the migrating group, enabling efficient collective migration colombi2021 .
@@
Proteomics, nucleic acids, and polymer modelling: Proteins are composed of amino acid chains and are vital for numerous biological functions. Their stability is influenced by factors such as temperature, pH, and interactions with other molecules. Cooperative self-assembly is a ubiquitous phenomenon found in natural systems dankulov2019 . However, these proteins lose their native structure through a process called “denaturation” due to environmental changes, which cause either reversible or irreversible changes in function. On the other hand, nucleic acids (DNA and RNA) are responsible for storing and transmitting genetic information. Biological polymers are large molecules composed of repeating structural units, and they are versatile, exhibit unique properties, and have various applications. Polymers for gene delivery systems occur across various length scales, from the molecular level to macroscopic structures, and aim to safely and effectively transport genetic material, such as DNA or RNA, into target cells paliy2009 ; paliy2010 ; badu2014 ; badu2020 ; badu2020a ; badu2020b . Polymeric fluids refer to liquids containing long-chain polymers and can exhibit complex rheological behaviour due to the interplay of molecular weight, concentration, and temperature yang2007c ; yang2008 ; melnik2002 ; melnik2003d ; melnik2003e . Furthermore, DNA damage within a biological network, such as a cellular or genomic network, can have profound implications for cellular function, integrity, and overall health.
@@
Dissipative processes in biosystems and structures: Dissipation is the conversion of energy from one form to another, which frequently results in the preservation of organized structures and the survival of life nicolis1971 . Living organisms are open systems that interchange matter and energy with their environment, allowing organized structures and functions to form. In addition, the non-equilibrium thermodynamic concepts regulate the dissipative processes in biosystems de2013 . The constant flow of energy maintains life by facilitating biochemical processes and sustaining dynamic equilibrium, such as the Belousov-Zhabotinsky reaction zhabotinsky2007 , from cellular metabolism to organismal homeostasis. These oscillations are examples of out-of-equilibrium dynamics that highlight biosystems’ ability to maintain order through energy dissipation.
@@
Synthetic biology and nonlocal model development: Synthetic biology is a multidisciplinary discipline that combines biological, engineering, and computer science ideas to produce synthetic creatures, genetic circuits, and biomolecular systems. Nonlocal models give a more complete picture of how biological components interact across distances inside a cell or a population of cells. This is especially important when creating synthetic biological systems with complex spatial and temporal relationships. These models are used in the construction of gene regulatory networks, allowing long-range interactions to be considered, which improves the accuracy of synthetic biology techniques for regulating gene expression hasty2000 . Furthermore, similar principles are extended to the study of population dynamics to include interactions between cells or animals that occur at a distance, which aids in the design of synthetic ecosystems and the understanding of emergent behaviours mee2012 . Nevertheless, wine fermentation involves the complex biochemical and physical processes transforming grapes into wine. The process is often characterized by the activities of numerous microorganisms, the consumption of sugars, and the creation of alcohol, carbon dioxide, and other byproducts. The dynamics of fermentation can be better understood using numerical modelling, particularly when it is based on integro-differential equations schenk2022 . As research in both domains advances, the incorporation of nonlocal models into synthetic biology adds to the development of novel biological tools, systems, and creatures with better spatial and temporal functions.
@@
Nonlocal models in systems biology: Systems biology seeks to understand the complex processes that occur inside biological systems by combining experimental data, computational modelling, and mathematical investigations. Nonlocal models are very useful in systems biology when dealing with distributed networks, such as signalling pathways, in which molecules can impact each other across distances inside a cell or tissue. It is also important to find the parameter values associated with the nonlocal model, which is called the inverse problem in systems biology oguz2013 . As collecting data and mathematical models advance, resolving challenges associated with inverse problems will lead to more precise and customized techniques in medicine, genetics, drug development, and other life science areas.
@@
Peridynamics: The application of peridynamics in biology is still an emerging area, and its adoption and development will rely on further advancements in computational techniques, experimental validation, and interdisciplinary collaborations between researchers in engineering, materials science, and biology taylor2016 ; zhou2022 . As the field progresses, peridynamics holds the potential to offer valuable insights into the mechanics and behaviour of biological systems under various conditions lejeune2021 .
@@
Biomimetics, bionics, and bio-inspired information processing: Genetic circuits and neuromorphic computation work in distinct biological environments, but their analogies emphasize the information processing, adaptability, and resilience principles that are critical to the operation of both systems. In addition, intercellular coupling is essential for multicellular organisms to function properly and maintain homeostasis. Disruptions in these communication pathways can lead to a variety of illnesses and developmental problems. The most significant type of natural computers may be found in biological systems that do computing on numerous levels chu2018 . Understanding these concepts can lead to breakthroughs in synthetic biology, biotechnology, and the creation of neuromorphic technological devices thieu2022 ; thieu2022a . Integration of signal and image processing with biology not only aids in the analysis of biological data but also in the development of computational models, diagnostic tools, and treatment methods. As bionic models evolve, driven by advances in materials, sensors, and robotics, the quality of life for people with disabilities will improve, human capacities will be enhanced, and numerous health concerns will be addressed ying2022 ; ying2022a ; ying2023 .

### v2

@@ 3.3 Physics, chemistry, and mechanics helping life science applications
The RD systems have also been applied to high energy physics (HEP) and statistical physics. RD models proposed by Balitsky and by Jalilian-Marian, Iancu, McLerran, Leonidov, Kovner, and Weigert (JIMLKW) have been used to understand the high energy hard scattering near or at the unitarity limit [442, 443]. Ikeda, Nagasawa, and Watanabe have introduced the branching Brownian motion in statistical physics [444, 445], which can be associated with the classical Fisher-KPP equation, as shown by McKean in [446]. Nonlocal interactions can occur due to long-range forces such as gravitational or electromagnetic interactions and can considerably impact the behaviour and dynamics of multi-body systems. These interactions are critical in astrophysics for analyzing the dynamics of galaxies because gravitational forces between stars and other celestial entities reach over large distances [447]. In galaxies, nonlocal interactions between stars and dark matter particles regulate galactic systems’ overall structure, rotation, and stability. Furthermore, the authors of [448] examined the self-interacting dark matter in cosmology, emphasizing the importance of nonlocal interactions in the dynamics of multi-body systems on cosmological scales.
Nonlocal variational formulations have been used to derive and study nonlocal PDEs, which appear in diverse physical and biological systems. For example, the nonlocal aggregation equation (50) has been adapted to describe biological organisms [449, 450, 451], and it minimizes energy
@@
and it can be used to analyze the existence and stability of equilibrium points [452, 453, 454]. Despite its popularity, it is not biological for a minimizer of mass $M$; doubling the mass yields a double-density minimizer. On the other hand, the dimensionless aggregation model (51) does form well-spaced groups. This model minimizes the energy [302, 455, 456]:
@@
These nonlocal variational problems are also used in image denoising and restoration tasks. The nonlocal operators capture the similarities between different image patches and help preserve important structures while reducing noise [457]. Furthermore, these problems are employed to study material’s deformation and fracture behaviour with long-range interactions and discontinuities, such as brittle materials and nanomaterials.
Nonlocal models also describe Ohmic heat production in different industrial processes, e.g., food sterilization, thermistor devices [458], etc. Using Ohmic heating, the nonlocal model of food sterilization in the one-dimensional spatial domain $(0,1)$ is given as [459]:
@@
with the initial condition $u(x,0)=u_{0}(x),0<x<1$ and the boundary condition $u(0,t)=u_{x}(1,t)=0,t>0$. Here, the function $f>0$ is the dimensionless electrical reactivity, and $\lambda>0$ is the square of the applied potential difference. A nonlocal hyperbolic model arising in Ohmic heating is examined in [460], whilst a nonlocal model with nonlinear diffusion is introduced and examined in [461, 462].
@@
Nonlocal models are used in many domains, including mechanics, materials science, and continuum physics. Nowadays, nanotechnology is one of the main technological interests in the research community, and nonlocal continuum mechanics could play a potential role in analyzing their applications [258, 463, 464]. Piezoelectric materials are used to make ultrasound scanners, and these sensors use the converse piezoelectric effect to send out sound vibrations reflected by muscles, organs, etc. [70]. These ultrasound devices also use the piezoelectric effect to detect the reflected sound. Then, they are displayed as images so that doctors can see those highly detailed 3D images for treatments. Other types of applications of these materials in the medical field are the making of sensors for electronic stethoscopes, hearing aids, flow meters to monitor vascular health and both external and internal sensors for health status monitoring. These materials are also used in the automobile industry, e.g., knock sensors, tyre pressure monitors, airbag and seatbelt sensors, gyroscopes, accelerometers, and engine fluid detection [465]. Piezoelectric and other smart materials and devices are increasingly important in the digital and physical world and human-machine interactions [466]. In these areas, nonlocal models, including degenerate, have proven advantageous [467, 468, 469, 470].
@@
Quantum dots are examples of a low-dimensional nanostructure, and researchers have been focused on these due to their several nanoelectronic applications [259, 471]. In the context of nano-inclusions and quantum dots, the size dependency of mechanical strain at the nanoscale has been brought to light in recent publications. Piezoelectricity is the growing importance of nonlocal models in the context of smart materials and structure [472]. Gradient theories are closer to microscopic theories like lattice dynamics than classical continuum theories. They are still applicable when the characteristic length of a problem is so small that classical continuum theories begin to fail [472]. The application of nonlocal continuum theory to nanotechnology was initially addressed by Peddieson et al. [258]. This model has been extended to the buckling and vibration analysis of multiwalled carbon nanotubes [62], which have been used extensively in biology and medicine for cancer treatment, drug delivery, and other therapeutic applications.
@@
Peridynamics is a nonlocal theory and is one of the key methods for modelling biological material fracture [275]. The main advantage of this nonlocal model is that it deals with continuous and discrete media [68]. It has been applied to find (i) the relationship between fracture and healing in cortical bone [66]; (ii) fracture in anisotropic cortical bone [273]; (iii) fracture in a complex wood microstructure [275]; (iv) rupture in a micron-scale biological membrane [67]; (v) fracture in a porous material [473]. Furthermore, peridynamics is used to model tissue growth and shrinkage. The distribution of division angle orientations influences tissue scale growth, and researchers have been trying to predict it in an individual cell [474]. Authors in [68] proposed a mechanics-based model with the help of the peridynamics framework. In addition, the authors have shown cell death differences on cellular and larger length scales. In particular, cell death may create gaps between cells in some cases, leading to tissue shrinkage in other cases.
@@
The nonlocal effects in heat-mass transfer arise in a wide variety of modern practical applications in materials sciences, e.g., metallic nanowires [475], high-power laser melting [476], colloidal solidification [477], frontal polymerization [478], etc. Many nonlocal mathematical and computational models developed for these and other applications [467, 469, 479, 480, 481, 482, 483] can be used as a basis for their extensions to nonlocal diffusion phenomena in a more general setting. The classical diffusion theory based on Fick’s law suggests the diffusion flux at a spatial point $x$, and a time moment $t$ depends on the concentration gradient at the same space-time point. These classical heat conduction laws apply at macro-temporal and spatial scales but not to thermal non-equilibrium situations, e.g., cryogenic engineering, laser-aided material processing, high-rate heat transfer in rarefied media, etc. Nowadays, researchers believe that heat-mass transport is an inherently nonlocal phenomenon. For instance, in one spatial dimension, a particle can come to a spatial point $x$ at a time moment $t+\tau$ from a spatial point $(x-l)$ or $(x+l)$ where it was at a time moment $t$. It implies a time lag $\tau$ between the diffusion flux and the particle concentration. This time lag $\tau$ is called relaxation time, particularly the mean collision time for gas particles. In the one-dimensional case, the net flux at a spatial point $x$ and time $t+\tau$ is given by
@@
After applying the zero-order approximation and Fick’s law, one can obtain the classical diffusion equation of the parabolic type. Furthermore, this approximation is valid for slow processes, i.e., $\tau\ll t_{0}$ and $l\ll h$, where $t_{0}$ and $h$ are the characteristic time and distances, respectively. In such cases, nonlocal effects are minimal, and the classical local equilibrium approach can be employed. For the other cases, e.g., $\tau\varpropto t_{0}$ and/or $l\varpropto h$, higher order approximation should be used as the nonlocal effects play a significant role. If the diffusion-like condition connects the small parameters:
@@
then, the first order approximation of (54) along with the mass conservation law $(\partial C/\partial t=-\partial J/\partial x)$, one can obtain [484, 485, 486]:
${}\frac{\partial C}{\partial t}+\tau\frac{\partial^{2}C}{\partial t^{2}}=d\frac{\partial^{2}C}{\partial x^{2}}+d\frac{l^{2}}{6}\frac{\partial^{4}C}{\partial x^{4}},$ (56)
where $d=\nu l=l^{2}/\tau$. This is a parabolic-type nonlocal equation, and for the same limiting condition (55), a higher-order approximation also gives the parabolic-type equation. For the wavelike dynamics, the limiting relation between the small parameters $\tau$ and $l$ is
@@
${}\frac{\partial C}{\partial t}+\tau\frac{\partial^{2}C}{\partial t^{2}}=d\frac{\partial^{2}C}{\partial x^{2}}.$ (58)
The limiting condition (57) also reduces to the hyperbolic type nonlocal equation for higher order approximation of (54). Therefore, the limiting conditions (55) and (57) mainly characterize the types of PDEs, whether it is parabolic or hyperbolic, and it is true for higher order approximations of (54).
The local nonequilibrium processes are described using extended irreversible thermodynamics, which incorporates higher-order fluxes into a set of independent variables [487]. The most basic evolution equation for the diffusion flux that accounts for both nonlocal and relaxation effects is:
$J+\tau\frac{\partial J}{\partial t}=-D\frac{\partial C}{\partial x}+l^{2}\frac{\partial^{2}J}{\partial x^{2}},$
where $l$ is the correlation length and $\tau$ is the relaxation time to local equilibrium [486]. Now, this equation with the mass conservation law, we obtain
${}\frac{\partial C}{\partial t}+\tau\frac{\partial^{2}C}{\partial t^{2}}=d\frac{\partial^{2}C}{\partial x^{2}}+l^{2}\frac{\partial^{3}C}{\partial t\partial x^{2}},$ (59)
which is of parabolic type and has an additional term (the mixed derivative) compared to the hyperbolic equation (58). Nevertheless, this equation coincides with the Jeffreys-type diffusion equation for liquids’ stress and strain rate [486, 487, 488].
Nonlocal models have received a lot of interest in the life and health sciences and provided an improved and comprehensive framework for analyzing a wide range of applications, including tissue growth, disease spread, neural networks, and tumor dynamics. The table below provides a summary of the mathematical models that have been covered in this section, emphasizing both their advantages and limitations.
Model Advantages Limitations
Extended Fisher-KPP model (46) It applies to a wider variety of systems, including ecosystems where competition is influenced by both local and nonlocal factors The exact analytical solutions are difficult to obtain, often requiring numerical methods and produce computational challenges
Age-structured models (48) and (49) These models incorporate the age-dependent infection, recovery, and mortality rates along with the long-range dispersal Getting high-resolution data on both age distribution and interaction patterns between different age groups
Model with conservation law (51) It conserves the zeroth and first moments of but not the higher moments This model neglects the birth, death, immigration, and emigration of organisms
Transport model (52) This model can integrate the chemotaxis for cell migration and follows the mass conservation law This model is limited to the condition in which the cues solely alter cell polarization, with a uniform distribution of speeds
Diffusion-convection model (53) This model can be applied to the food industry, illustrating the evolution of the temperature of the sterilized food Different solution approaches should be followed, though depending on the monotonicity of the nonlinearity because no maximum principle is available for such nonlocal parabolic problems
Nonlocal diffusion models (56), (58), and (59) These models are based on the ratio of the mean free path to the characteristic length of the system and can be used for the model of the processes far from equilibrium conditions These approaches can not be applied to a system if it does not follow such specific ratio
@@ 4 Emerging Trends, Future Directions, and Outlook
The areas of mathematical modelling in biology are still evolving; various developing patterns in nonlocal mathematical models have been observed. As research progresses, these models are expected to play an increasingly significant role in addressing new challenges in biology and life sciences. These trends reflect advances in computational approaches, enhanced experimental data availability, and a rising acknowledgement of the importance of nonlocal interactions in understanding complex biological phenomena. Some of the emerging trends in nonlocal mathematical models are the following.
@@
Multi-scale modelling: Biological systems often span multiple spatial and temporal scales, and nonlocal models provide a robust and powerful framework to bridge these scales and capture the intricate interactions between different components of the system, from molecular to cellular to tissue and organism levels, however, traditional modelling approaches often focus only on specific scales [64, 65].
@@
Human-environment systems, biodiversity, and ecological modelling: Human-environment systems represent the intricate relationship between human societies and their surrounding environments [489]. In addition, biodiversity is the variety of life on Earth, encompassing different species, ecosystems, and genetic diversity. It is critical for maintaining ecological balance, delivering ecosystem services, and contributing to ecosystem resilience in the face of environmental change. Ecological modelling is developing and analyzing mathematical and computer models to replicate and comprehend the dynamics of ecosystems. Studying the spatio-temporal trends of human activities, such as deforestation rates, carbon emissions, and resource consumption, provides insights into the evolving nature of anthropogenic impacts on the environment. These models assist researchers in predicting the impact of human activities on biodiversity, studying ecosystem responses, and developing sustainable resource management techniques.
@@
Collective cell migration: Collective cell migration is a phenomenon where groups of cells move together to achieve specific functions, such as embryonic development, wound healing, or cancer invasion [75, 76, 52, 490]. These nonlocal interactions can influence cell behaviour, signalling, and collective dynamics in complex biological systems. These interactions play a significant role in coordinating the movements of cells within the migrating group, enabling efficient collective migration [491].
@@
Proteomics, nucleic acids, and polymer modelling: Proteins are composed of amino acid chains and are vital for numerous biological functions. Their stability is influenced by factors such as temperature, pH, and interactions with other molecules. Cooperative self-assembly is a ubiquitous phenomenon found in natural systems [492]. However, these proteins lose their native structure through a process called “denaturation” due to environmental changes, which cause either reversible or irreversible changes in function. On the other hand, nucleic acids (DNA and RNA) are responsible for storing and transmitting genetic information. Biological polymers are large molecules composed of repeating structural units, and they are versatile, exhibit unique properties, and have various applications. Polymers for gene delivery systems occur across different length scales, from the molecular level to macroscopic structures, and aim to safely and effectively transport genetic material, such as DNA or RNA, into target cells [493, 494, 495, 496, 497, 498]. Polymeric fluids refer to liquids containing long-chain polymers and can exhibit complex rheological behaviour due to the interplay of molecular weight, concentration, and temperature [499, 500, 501, 502, 503]. Furthermore, DNA damage within a biological network, such as a cellular or genomic network, can have profound implications for cellular function, integrity, and overall health.
@@
Dissipative processes in biosystems and structures: Dissipation is the conversion of energy from one form to another, which frequently results in the preservation of organized structures and the survival of life [504]. Living organisms are open systems that interchange matter and energy with their environment, allowing organized structures and functions to form. In addition, the non-equilibrium thermodynamic concepts regulate the dissipative processes in biosystems [289]. The constant flow of energy maintains life by facilitating biochemical processes and sustaining dynamic equilibrium, such as the Belousov-Zhabotinsky reaction [505], from cellular metabolism to organismal homeostasis. These oscillations are examples of out-of-equilibrium dynamics highlighting biosystems’ ability to maintain order through energy dissipation.
@@
Synthetic biology and nonlocal model development: Synthetic biology is a multidisciplinary discipline that combines biological, engineering, and computer science ideas to produce synthetic creatures, genetic circuits, and biomolecular systems. Nonlocal models give a more complete picture of how biological components interact across distances inside a cell or a population of cells. This is especially important when creating synthetic biological systems with complex spatial and temporal relationships. These models are used in constructing gene regulatory networks, allowing long-range interactions to be considered, which improves the accuracy of synthetic biology techniques for regulating gene expression [506]. Furthermore, similar principles are extended to the study of population dynamics to include interactions between cells or animals that occur at a distance, which aids in designing synthetic ecosystems and understanding emergent behaviours [507]. Nevertheless, wine fermentation involves the complex biochemical and physical processes transforming grapes into wine. The process is often characterized by the activities of numerous microorganisms, the consumption of sugars, and the creation of alcohol, carbon dioxide, and other byproducts. Fermentation dynamics can be better understood using numerical modelling, particularly when it is based on integro-differential equations [508]. As research in both domains advances, incorporating nonlocal models into synthetic biology adds to developing novel biological tools, systems, and creatures with better spatial and temporal functions.
@@
Nonlocal models in systems biology: Systems biology seeks to understand the complex processes inside biological systems by combining experimental data, computational modelling, and mathematical investigations. Nonlocal models are very useful in systems biology when dealing with distributed networks, such as signalling pathways, in which molecules can impact each other across distances inside a cell or tissue. It is also important to find the parameter values associated with the nonlocal model, called the inverse problem in systems biology [509]. As collecting data and mathematical models advance, resolving challenges associated with inverse problems will lead to more precise and customized techniques in medicine, genetics, drug development, and other life science areas.
@@
Peridynamics: The application of peridynamics in biology is still an emerging area, and its adoption and development will rely on further advancements in computational techniques, experimental validation, and interdisciplinary collaborations between researchers in engineering, materials science, and biology [67, 510]. As the field progresses, peridynamics holds the potential to offer valuable insights into the mechanics and behaviour of biological systems under various conditions [68].
@@
Biomimetics, bionics, and bio-inspired information processing: Genetic circuits and neuromorphic computation work in distinct biological environments, but their analogies emphasize that information processing, adaptability, and resilience principles are critical to operating both systems. In addition, intercellular coupling is essential for multicellular organisms to function properly and maintain homeostasis. Disruptions in these communication pathways can lead to various illnesses and developmental problems. The most significant type of natural computers may be found in biological systems that do computing on numerous levels [511]. Understanding these concepts can lead to breakthroughs in synthetic biology, biotechnology, and the creation of neuromorphic technological devices [512, 513]. Integration of signal and image processing with biology not only aids in analysing biological data but also in developing computational models, diagnostic tools, and treatment methods. As bionic models evolve, driven by advances in materials, sensors, and robotics, the quality of life for people with disabilities will improve, human capacities will be enhanced, and numerous health concerns will be addressed [514, 515, 516].